# Privacy-Preserving Dataset Combination

### v2

@@ 3.3 KL-based Methods, in Plaintext
Specifically, $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ score used in $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ first trains a logistic regression model [37] on $\mathcal{D}_{o}\cup\mathcal{D}_{i}$ – where the labels are folded into the covaraiates – with the goal of inferring dataset membership. A score of $0.5$ or less means the datasets are not distinguishable, making the data potentially useful.  [13] established the insight that in data-limited domains of heterogeneous data sources, domain shifts of the covariates are useful for predicting whether the additional data helps the original task. We note again that even though this model is trained on both parties’ data, the final algorithm that the hospital uses to train on combined data is not restricted.
@@
We focus on $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$, and reproduce [13] ’s results that it is predictive of downstream change in AUC.
Let the score function $g_{\mathrm{KL}}$ be the approximate of $\mathrm{KL}(\mathcal{D}_{o}||\mathcal{D}_{i})$. The strategy selects the most likely hospital with the closest distance under the measure:
@@ 3.4 SecureKL: Private KL-based Method
Using MPC, we extend on $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to require no information sharing (besides what was already assumed public). Specifically we leverage the MPC based framework provided by CrypTen [35], a library designed for privacy-preserving machine learning, to implement private $\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$. As illustrated in Figure 1, the logistic regression as well as the scoring need to be implemented in private. Our code is publicly available  [^1] https://anonymous.4open.science/r/Private-Preserving_Data_Combination-451E.
@@
As shown in Figure 1, any KL-based measure $g_{\mathrm{SKL}}$ can be adapted to our setup. We mainly use $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ as the underlying measure. Its performance is detailed in Section 4.4. Additionally, even though our implementation measures distance of data between one source and one target party, the setup readily extends to accommodating multiple parties. We note the engineering limitations in Section 4.5.3.
@@ 4 Experiments
4.1 Experimental Setup
We validate our method and demonstrate its applicability using the eICU Collaborative Research Dataset [38], which contains over $200,000$ admissions from $208$ hospitals across the United States. Following the data cleaning and exclusion criteria outlined by [39] and [13], we selected the $12$ hospitals with the highest number of patient visits (each with at least $2000$ patients) as our $\mathbf{H}$. Each strategy would compute with the same $K=3000$ records, as the total available data per hospital.
We simulate the problem setup for each hospital with the $24$-hour mortality prediction task. The strategy comparisons described in Section 3 are implemented using $1500$ samples and the AUC is evaluated on $400$ samples for all of our experiments unless otherwise noted. This follows training and evaluation protocols in Yet Another ICU Benchmark [39]. For the data combination experiments that compute AUC change $\delta_{i}$ or $\delta_{T}$, to match  [13] , we take $1500$ random samples from each selected dataset and combine it with $1500$ samples from $\mathcal{D}_{o}$ (fixed across all experiments).
Implementing $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be privacy-preserving requires training logistic regression model in private. This is used in the private setting to estimate $g_{\mathrm{SKL}}$$\mathrm{Score}([X])$ or $\mathrm{Score}([X,Y])$ – for each pair of hospitals. Our experiments train encrypted logistic regression in CrypTen [35] using the library’s SGD optimizer. To ensure a fair comparison between the scores obtained through plaintext and encrypted settings, we re-implement plaintext Score($X$) and Score($X,Y$) using logistic regression with SGD in PyTorch [40]. This is because encrypted version of L-BFGS – the optimizer prior work [13] uses in plaintext-only with Scikit-Learn [41]– is not available in CrypTen, though it leads to better downstream performance. Hyper-parameter tuning for SGD in private and plain text are performed independently, with the details in Appendix 8.
@@ 4.2 Experimental Questions
    Consistency: Does using multiparty implementation sacrifice original measure’s effectiveness? Practically, we evaluate this through the analysis of private and plaintext scores. For our selected metrics, we expect AUC change to be negatively with KL-based measures—meaning the closer the additional target dataset is to the source hospital, the more the AUC will improve compared to other potential target datasets. Section 4.3 tests the correlation of our private scores and plaintext scores with full access (setting $k=K$). In addition to computing Spearman’s rank correlation coefficient of the KL-scores, we probe the discrepancy of the downstream effect between $\{\delta_{i}|i=\pi_{s}(n=1,k=K)\}$ using plaintext $\mathrm{KL}_{\mathcal{X}},\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ with all the underlying data and $\{\delta_{i}|\pi=\pi_{p}\}$ using $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}},\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ for each source hospital $H_{o}\in\mathbf{H}$.
@@
    Positivity: Does our method pick hospitals that reliably improve performance? If source dataset $D_{o}$ can only add data from $n$ more hospitals, does our measure lead to eventual AUC improvements? In Section 4.4, we test our framework on a multi-dataset combination experiment and find that it successful improves the source hospital’s downstream outcome. Specifically, when selecting a single additional data source ($n=1$), all but 2 hospitals improves, and when selecting top 2 or 3, all hospitals see a positive AUC change. This shows a consistent added benefit.
    Lastly, Section 4.4.1 compare with different strategies proposed in Section 3, and Section 4.4.2 analyze the benefits of using private dataset combination $\mathrm{SKL}$.
@@
    Error analysis: If our privacy-preserving method is not the dominant strategy against alternatives including limited data accessibility, why is that? Section 4.5 performs additional analyses on (a) hospitals with low $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ correlations, and (b) hospitals lagging AUC improvements using the random strategy $\pi_{0}$ or limited-sample strategy $\pi_{s}$ for selecting $n=3$ candidate hospitals.
Lastly, we note engineering hurdles to scale to the real-world in Section 4.5.3. We hereby detail our results.
@@ 4.3 Consistency Between Plaintext and Encrypted Computations
Our encrypted computations are programmed with CrypTen, when the plaintext counterpart is PyTorch-only. We show using $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ lead to highly comparable behavior as $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{KL}_{\mathcal{X}}$, respectively.
@@
For each source hospital $H_{o}$, use all full samples for $\mathcal{D}_{i}$. Between $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ on $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ for all remaining hospitals $H_{i}$, $\mathbb{E}_{H_{o}\sim\mathbf{H}}[\rho]=0.908$ with a range of $[0.691,1.0]$, obtaining $p<0.02$ across all hospitals. Between $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}}$, $\mathbb{E}_{H_{o}\sim\mathbf{H}}[\rho]=0.9303$ with a range of $[0.455,0.991]$, with 11 of 12 hospitals achieving p-values below $0.05$. After applying Hochberg false discovery rate correction [42], our p-values remain significant. This range is an artifact of sweeping hyperparameters independently in plaintext and encrypted optimisations. When we unify the SGD hyperparameters, we indeed get a tighter range. For all 12 hospitals, see appended Appendix 9 for details.
Downstream Model Improvements
@@
We further simulate the effect by adding encryption through its impact on the downstream AUC. This examines whether there will be a shift in the full hospital ranking, if we switch from a plaintext setup to encrypted. For $H_{o}\sim\mathbf{H}$, we measure $\delta_{i}$ that results from adding $\mathcal{D}_{i}$ to $\mathcal{D}_{o}$ for all $i$. This correlates all target hospitals $\{H_{i}\}$ with their ground truths $\{\delta_{i}\}$ in the case of picking a single target hospital. We find the linear coefficient for encrypted $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be $-0.182$ and plaintext $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be $-0.184$ ($99\%$ matching). Both $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}}$ have a linear coefficient of $-0.164$ with $\delta_{i}$. For all strategies’ correlations with ground truth at $n=1$, see Appendix 7.
Figure 2: AUC change $\delta_{T}$, after including top-$3$ hospital per strategy, over all source hospitals. Our private strategy $\pi_{p}$ is compared with (a) demographic-based $\pi_{d}$ for gender, age, and race. (b) plain text limited-sample $\pi_{s}(k=K)$ for $k=300$ (10%) and $k=30$ (1%).
4.4 Positivity of SecureKL and Practical Implications
Overall Positivity
@@ 4.4 Positivity of SecureKL and Practical Implications
We evaluate the practical utility of SecureKL by applying it in a multi-source data combination experiment, where $n\in\{1,2,3\}$. For $n=1$, we find that $\pi_{p}$ improves AUC in 10 out of the 12 hospitals. When $n=2$ and $n=3$, we find that using $\pi_{p}$ consistently improves AUC for all hospitals. Overall, 34 out of the 36 dataset combinations we evaluate on have an AUC improvement $\delta_{T}>0$, suggesting that $\pi_{p}$ is a reasonable strategy for selecting hospital dataset combinations with a high expected return $\mathbb{E}[P_{H_{o}\sim\mathbf{H}}(\delta_{T}>0)]$ for the source hospital from using our strategy.
4.4.1 Comparing With Alternative Strategies
@@
Other strategies – $\pi_{0}$, $\pi_{d}$, and $\pi_{s}$ – can also arrive at positive datasets. Comparing private method $\pi_{p}(n=3)$ to other strategies at $n=3$, we find the following results, illustrated in Figure 2:
@@
[…]
@@
In summary, our method $\pi_{p}$ achieves the highest AUC improvement on average with the lowest standard deviation, demonstrating consistent improvement for all hospitals. In contrast, demographic-based and plaintext small-sample strategies exhibit greater variability, with some strategies improving performance for specific subsets of hospitals but underperforming or degrading results in others.
4.4.2 SecureKL Analysis
After establishing that $\pi_{p}$ with $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$is a robust strategy in practical downstream performance, we hereby synopsize the benefits of $\mathrm{SecureKL}$ and elaborate on their practical implications.
A Principled Approach To Data Minimization. Our major contribution is to match plaintext performance with no data sharing. Using MPC provides input privacy, meaning that if both hospitals only want to know the resulting score, the computation can be done without leaking original data. This strong guarantee can significantly ease the tension related to privacy and compliance in setting up a collaboration, leading to a practical ”data appraisal stage” in data-limited high stakes domains.
In the case where that output can be sensitive, i.e., when hospitals query each other multiple times and accrue information through the score function, the output can also be made privacy-preserving through differentially private data releases, such as using randomized response [43].
In theory, any data combination method (if Turing-complete) can be made private; yet, in practice, balancing the right trade-off of utility and privacy is non-trivial. Barring engineering difficulties, not all algorithms readily adapt efficiently in private. Prior work  [44] included the trained model and test data in private; while relatively exact, complex methods would exacerbate the same operational limitations discussed in Section 4.5.3.
Gain from Data Availability. In contrast to limited-sample approaches, a key advantage for our method $\pi_{p}$ is that it takes advantage of all of the underlying data – generally impossible with non-secure methods for private data in heavily regulated domains. The general intuition is that data is localized; therefore, once a good target hospital is identified, we should acquire all of the data. It may be tempting to assert that we prefer the highest $k$ for data addition algorithms as well. In our experiments, while this is generally true, the smaller $k$ sometimes outperform larger $k$ in plaintext strategy $\pi_{s}$, which we investigate in Section 4.5 and in Figure 5. This occasionally non-monotonic behavior mirrors the challenge of data combination itself: even within one source dataset for the same estimator, more data is not necessarily better. This suggests a domain-specific alternative to sharing a large amount of data for some source hospital, and points to future directions to using secure computation on a minimal-sized sample dataset for minimal performance overhead while remaining private.
4.5 Error Analysis
4.5.1 Underlying Score Limitations

### v3

@@ 4.2 KL-based Methods, Without Privacy
Specifically, $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ score used in $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ first trains a logistic regression model (Cox, 1958) on $\mathcal{D}_{o}\cup\mathcal{D}_{i}$ – where the labels are folded into the covariates. The goal is to infer dataset membership using “proxy labels”, defined as follows:
$\mathcal{I}(x,y)=\begin{cases}1&\text{if }(x,y)\in\mathcal{D}_{o}\\ 0&\text{if }(x,y)\in\mathcal{D}_{i}\end{cases}$
A binary predictor is fit on this combined dataset, predicting $\mathcal{I}$ from $(\mathcal{X},\mathcal{Y})$ using logistic regression. The model’s output $p(x,y)$, also called the probability score, is Score(x,y)(x,y).
A score of $0.5$ or less means the datasets are not distinguishable, making the data potentially useful. Shen et al. (2024) established the insight that in data-limited domains of heterogeneous data sources, domain shifts of the covariates are useful for predicting whether the additional data helps the original task, similar to Miller et al. (2021). We note again that even though this model is trained on both parties’ data, the final algorithm that the hospital uses to train on combined data is not restricted.
@@
We reproduce Shen et al. (2024) ’s results that $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ is predictive of downstream change in AUC in hospital setting (Table 4). Let the score function $g_{\mathrm{KL}}$ be the approximate of $\mathrm{KL}(\mathcal{D}_{o}||\mathcal{D}_{i})$. The strategy selects the most likely hospital with the closest distance under the measure:
@@
Though leakage can be controlled through $k$, yet the data is inherently sensitive. In ICU data, simulate that a default of $1\%$ is shared, so $k=3000\times 1\%=30$, though we run experiments with $k\in\{3,30,300,3000\}$.
@@ 4.3 SecureKL: Private KL-based Method
Figure 4: Our method SecureKL. Each side encrypts their data. Then, a model is privately trained on their joint data. Afterwards, their divergence is computed. Finally, only the final result of this dataset-to-dataset evaluation is revealed.
Using MPC, we extend on $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to require no information sharing (besides what was already assumed public). Specifically we leverage private tensors and secure gradient descent in CrypTen (Knott et al., 2021) to implement private $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$. As illustrated in Figure 4, the logistic regression as well as the scoring need to be implemented in private.
@@
As shown in Figure 4, any KL-based measure $g_{\mathrm{SKL}}$ can be adapted to our setup. We mainly use $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ as the underlying measure. Its performance is detailed in Section 6.2. Additionally, even though our implementation measures distance of data between one source and one target party, the setup readily extends to accommodating multiple parties. Section 7.2 discusses potential deployment challenges.
Figure 5: SecureKL: Overall Correctness. Rank correlation between SecureKL output and ground truth AUC change, $\delta_{i}$, from acquiring 11 additional dataset for a given source hospital $H_{o}$. We propose selecting data partner ranked by our secure system under $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ score to reliably increase AUC gains. ($|\mathbf{H}|=12$ hospitals; colored by source.)
4.4 Trivially “Private” Baseline: Blind Selection
$\pi_{0}(n)$ randomly selects $n$ disjoint data sources, until data purchasing budget runs out. This random strategy may evade selection biases and help gather diverse data. Yet, prior work  (Shen et al., 2024) suggests that $\pi_{0}(1)$ – randomly selecting one source – for ICU is risky and inefficient for mortality prediction.
4.5 Alternative Data-leaking Baseline: Sharing Summary Statistics
@@ 5.1 Experimental Questions
    Consistency: Does MPC degrade the original measure’s effectiveness? Since MPC implementations introduce approximations, we first validate $\mathrm{SKL}$’s correctness by examining the correlation of private scores (SecureKL) vs. plaintext scores ($\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$, with full data access). Then, for the consistency of choosing encryption in the hospital domains, we examine plaintext and encrypted versions’ correlation with downstream ground truth rankings across hospitals.
@@
    Positivity: Does our method pick entities that reliably improve performance? If source dataset $D_{o}$ can only add data from $n$ more sources, does our measure lead to eventual AUC improvements? Specifically, in multi-dataset combination, we examine whether using $\mathrm{SKL}$ can improve the source hospital’s downstream outcome. When selecting a single (or a few) additional data source, how many hospitals improve with our method? Additionally, we compare our method with alternative, privacy-leaking strategies.
@@
    Error analysis: If our privacy-preserving method is not the dominant strategy against alternatives including limited data accessibility, why? As we know, small and uncertain improvements for downstream tasks underscore the inherent difficulty of evaluating data utility without seeing the full data. We are especially interested in analyzing (a) hospitals with low $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ correlations, and (b) hospitals lagging AUC improvements using the random strategy $\pi_{0}$ or limited-sample strategy $\pi_{s}$.
5.2 MPC implementation
$\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ includes training logistic regression model in private. We implement measures based on dataset divergence by building a custom logistic regression model over encrypted data leveraging CrypTen. The model parameters and input are encoded as 16-bit MPCTensors, ensuring that all computations, including forward passes, sigmoid activations, and gradient descent updates, are performed in private.
Additional baseline details
We additionally run our experiments on plaintext methods used in Shen et al. (2024), including the $\mathrm{KL}_{\mathcal{X}}$ measure, which is similar to $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ without using each data source’s labels:
To compare against $\mathrm{KL}_{\mathcal{X}}$ for baseline, we additionally implemented its encrypted version, $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$, as a $g_{\mathrm{SKL}}$ candidate.
Optimizers
Because L-BFGS – the optimizer prior work (Shen et al., 2024) used in plaintext-only with Scikit-Learn (Pedregosa et al., 2011) – is not available as an encrypted version, our MPC experiments are facilitated with SGD optimizer. A fair comparison between the scores obtained through plaintext and encrypted settings necessitates re-implementing plaintext scores, Score($X$) and Score($X,Y$), using logistic regression with SGD in PyTorch (Paszke et al., 2019). The hyper-parameter tuning for SGD in private and plain text are performed independently, as they do not transfer. Appendix 12 will discuss hyperparameter-tuning in detail.
5.3 Data and Model Setup
eICU dataset. The downstream task is $24$-hour mortality prediction from ICU data using the eICU Collaborative Research Dataset (Pollard et al., 2018). This dataset contains >200,000>200,000 real-world admission records from $208$ hospitals across the United States.
A note on data filtering. Because medical research is inherently complex, ensuring reproducibility on statistical methods can be significantly challenging.  Water et al. (2023) proposes that the research community follow a shared set of tasks with fixed preprocessing pipelines that are clinically informed – essentially a machine learning training and evaluation protocol on eICU – in order to facilitate method verification on the same benchmark. Therefore, our work follows their data cleaning criteria and the evaluation protocol. Additionally, we use the hospital exclusion criteria in (Shen et al., 2024) to obtain top $12$ hospitals, where the most patient visits are collected (each with >2000>2000 patients).
Downstream model and baselines for eICU. Recall that each strategy uses the same $K$ number of records per hospital – in our experiment, $K=3000$. For $\pi_{s}$ which leaks a subset $k$ of all samples, a default $k=1\%|\mathcal{D}_{i}|$ randomly drawn samples are shared. In ICU data, we run experiments on $\{0.1\%,1\%,10\%,100\%\}$.
Following holistic benchmarking tools in  (Water et al., 2023), our strategy comparisons take $1500$ samples and the downstream model performance – $\mathrm{AUC}_{o}$, $\mathrm{AUC}_{T}$ – uses $400$ samples (unless otherwise noted). Specifically, the AUC change, $\delta_{i}$ or $\delta_{T}$, comes from 1. combining $1500$ random samples from each selected dataset and 2. combining it with $1500$ samples from $\mathcal{D}_{o}$, and 3. subtracting the baseline model’s AUC [^2] The samples are fixed across all experiments, the sample numbers are chosen to match (Shen et al., 2024)’s setup..
Folktables dataset Though we primarily focus on hospital domain, we additionally validate using Folktables (Ding et al., 2021), predicting across 15 states an individual’s annual income exceeds \$50,000333Because the Folktables dataset contains simpler features and is low-dimensional, KL divergence is directly computed rather than estimated. For implementation details, see Appendix 15.. The details of our processing, which diverges from that of eICU, is included in Appendix 15.
6 Results and Analysis
We organize our results around three research questions: whether using multiparty implementation sacrifices $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$’s efficacy (consistency; Section 6.1), whether our method reliably picks hospitals that improve performance (positivity; Section 6.2), and where our method may fail (error analysis and limitations; Section 6.3).
@@ 6.1 Consistency Between Plaintext and Encrypted Computations
Because our encrypted computation is the first implementation of dataset divergence in MPC, we ought to show that $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ lead to highly comparable behaviors as $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{KL}_{\mathcal{X}}$.
@@
For each source hospital $H_{o}$, use all full samples for $\mathcal{D}_{i}$. Between $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ on $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ for all remaining hospitals $H_{i}$, $\mathbb{E}_{H_{o}\sim\mathbf{H}}[\rho]=0.908$ with a range of $[0.691,1.0]$, obtaining $p<0.02$ across all hospitals. Between $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}}$, $\mathbb{E}_{H_{o}\sim\mathbf{H}}[\rho]=0.9303$ with a range of $[0.455,0.991]$, with 11 of 12 hospitals achieving p-values below $0.05$. After applying Hochberg false discovery rate correction (Benjamini and Hochberg, 1995), our p-values remain significant. This range may be an artifact of sweeping hyperparameters independently in plaintext and encrypted optimisations, because sharing the same SGD hyperparameters would result in a tighter range. For all 12 hospitals, see appended Appendix 13 for details.
Impact of Adding Security on AUC Correlation
@@
We further examine the effect by adding encryption through its impact on the downstream AUC, using how AUC improvements are ranked. This rank is compared with how secure measures (i.e., $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$) and plaintext measures (i.e. $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$) rank hospitals. This comparison investigates the extent of the shift in the full hospital ranking, when we switch from a plaintext setup to encrypted. For $H_{o}\sim\mathbf{H}$, we measure $\delta_{i}$ that results from adding $\mathcal{D}_{i}$ to $\mathcal{D}_{o}$ for all $i$. This correlates all target hospitals $\{H_{i}\}$ with their ground truths $\{\delta_{i}\}$ in the case of picking a single target hospital. We find the linear coefficient for encrypted $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be 0.182-0.182 and plaintext $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be 0.184-0.184 ($99\%$ matching). Both $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}}$ have a linear coefficient of 0.164-0.164 with $\delta_{i}$. For all strategies’ correlations with ground truth at $n=1$, see Appendix 11.
6.2 Positivity in Realistic Setup
We apply SecureKL in a pragmatic multi-source data combination problem, where each strategy acquires $n$ datasets for $n\in\{1,2,3\}$.
Overall Positivity.
@@ 6.2 Positivity in Realistic Setup
For $n=1$, we find that $\pi_{p}$ improves AUC in 10 out of the 12 hospitals. When $n=2$ and $n=3$, we find that using $\pi_{p}$ consistently improves AUC for all hospitals. Overall, 34 out of the 36 dataset combinations we evaluate on have an AUC improvement $\delta_{T}>0$, suggesting that $\pi_{p}$ is a reasonable strategy for selecting hospital dataset combinations with a high expected return $\mathbb{E}[P_{H_{o}\sim\mathbf{H}}(\delta_{T}>0)]$ for the source hospital from using our strategy.
Demographic Summary Strategies Subset-sampling Strategies Secure Strategy (Ours)
Dataset $n$ $\pi_{d}$-gender $\pi_{d}$-age $\pi_{d}$-race $\pi_{s}(k=300)$ $\pi_{s}(k=30)$ $\pi_{p}\ \mathrm{Secure}\mathrm{KL}_{\mathcal{XY}}$
eICU 1 $\textbf{0.020}\pm 0.023$ $012\pm 0.016 0.0$$14\pm 0.015 0.0$$12\pm 0.014 0.0$$\pm 0.017 0.011±0.0$$1\pm 0.018$
2 $016\pm 0.016 0.0$$17\pm 0.017 0.0$$15\pm 0.013 0.0$$24\pm 0.020 0.0$$17\pm 0.019 0.0$$2\textbf{0.027}\pm$ 0.022
[…]
2 $\textbf{0.015}\pm 0.019$ $\textbf{0.015}\pm 0.014$ $0\pm 0.013 0.011±0.0$$1\pm 0.014$ $014\pm 0.018 0.0$$010\pm 0.017$
[…]
Table 2: AUC improvements in mean and standard deviation, across all source regions for each strategy $\pi$, for eICU and Folktables setups. $n$ denotes the number of candidate datasets added to the source dataset. The small gains and high variance from adding selected datasets highlight the precarious nature of assessing data value in the real world. Bold indicates highest AUC improvement per $n$. Note: Only $\pi_{p}\ \mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ is private.
Figure 6: AUC change $\delta_{T}$ over all strategies in eICU prediction (higher is better). Our private dataset evaluation strategy $\pi_{p}$ outperforms demographic-based strategy $\pi_{d}$ (left), and subset-sampling strategy $\pi_{s}$ for $k=300$ (10%) and $k=30$ (1%) (right), after combining source data with the top 33 candidates.
Figure 7: AUC change $\delta_{T}$ over all strategies in Folktables dataset prediction (higher is better). All strategies exhibit comparable distributions, after combining data from top 3 candidates. In a noisy domain, our method is stable: it neither excels nor penalizes against non-private strategies.
Comparing With Alternative Strategies
@@
Other strategies – $\pi_{0}$, $\pi_{d}$, and $\pi_{s}$ – can also arrive at “positive” datasets. Comparing private method to other strategies at $n=3$, i.e, $\pi_{p}(n=3)$, we describe our results in Figure 7:
@@
    $\pi_{p}$ (our method based on $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$) has a median $\delta_{T}$ of $0.020$, and a standard deviation of $0.015$. Our results indicate that for 5050% of the hospitals, $\pi_{p}$ gives a $\delta_{T}>=.02$. Compared to other strategies, $\pi_{p}$ has the highest median, the lowest standard deviation, and it is one of two strategies that improves performance for all hospitals.
@@
In summary, our method $\pi_{p}$ achieves the highest AUC improvement on average with the lowest standard deviation, demonstrating more consistent improvements for all hospitals. While the average improvement of $\pi_{p}$ is small, demographic-based and plaintext small-sample strategies exhibit greater variability, with some strategies improving performance for specific subsets of hospitals but underperforming or degrading results in others.
6.3 SecureKL Error Analysis
Underlying Data Limitations
In high-stakes domains, data partnerships are expensive, but potentially detrimental – this forms a challenging landscape for evaluating methods on real-world data. Indeed, as shown in Table 2, the AUC gain is small across all strategies, and the variance is high. This suggests that 3 hospitals’ data is likely still too small for the general task to the robust explains limited AUC gains, highlighting the need to maximize samples for informative decisions. The key distinction, however, is that privacy-leaking methods (demographic, small sample) and blind baseline risk performance declines in many hospitals while SKL consistently improves downstream tasks more reliably than alternatives, across all hospitals, over multiple runs.
Figure 8: The correlation between AUC change $\delta_{i}$ and underlying score $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ affects $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$’s efficacy. In Hospital 420, the $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ scores identify beneficial data ($\rho<0$). For Hospital 243, $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ anti-correlates ($\rho>0$).
Underlying Score Limitations